Nesmith Corporation's outstanding bonds have a $1,000 par value, a 12% semiannual coupon, 16 years to maturity, and a 7% YTM. What is the bond's price? Round your answer to the nearest cent.
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, n = 16 * 2 = 32 semi-annual periods, i = 7%/2 = 3.5% (semi-annual), C = 12% * $1000/2 = $60

1- (1+0.035)52 P = 60 * - 0.035 - - - 1000 (1 + 0.035 32

P = $1,144.1 + $332.6

P = $1,476.7